Solve x2 + 2x = 4 for x by completing the square. Ox=+15+1 Ox=315-1 Ox=1 x = 3
Answer #1

Hey there stay home and stay safe

\small x^2+2x=4\\ \Rightarrow x^2+2\times x\times 1=4\\ \Rightarrow x^2+2\times x\times 1+1=4+1\\ \Rightarrow x^2+2\times x\times 1+1^2=5\\ \Rightarrow (x+1)^2=5\\ \Rightarrow x+1=\pm\sqrt5\\ \Rightarrow x=-1+\pm\sqrt5\\

And please leave a like. Thank you.